Find the mean of the probability distribution of a random variable X which can take 

only the values 3, 5, and 7, given that P(3)=7/30, P(5)= 1/3, and P(7)=13/30


1
Expert's answer
2022-03-23T02:02:30-0400

"\mu=3*\frac{7}{30}+5*\frac{1}{3}+7*\frac{13}{30}=\frac{162}{30}=5.4."